OpenStudy (anonymous):

could someone please explain to me how to add or subtract rational expressions

OpenStudy (anonymous):

well a/b + c/d = (a x d + b x c)/ b x d

OpenStudy (anonymous):

and a/b - c/d = (a x d - b x c)/ b x d

OpenStudy (anonymous):

here x = multiplication

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

if the base is same you simply add or subtract upper part keeping the base same

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you aajugdar, maybe I am just thinking to deep into this

OpenStudy (anonymous):

its easy and basic dear. :)

OpenStudy (anonymous):

right brain...left brain I guess ;)

OpenStudy (anonymous):

heheh well did u get the method?

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

good xP

OpenStudy (anonymous):

here is an example to keep me busy \[\frac{x}{x-3}+\frac{2}{x+1}=\frac{x(x+1)+2(x-3)}{(x-3)(x+1)}\]